OpenStudy (samsan9):

find the vertex, focus, directrix, and the length of the focal chord for y^2+6y-2x+13=0

ganeshie8 (ganeshie8):

change it to vertex form

OpenStudy (samsan9):

how do you change it ?

OpenStudy (samsan9):

y=-13(x-2)?

ganeshie8 (ganeshie8):

\[\large y^2+6y-2x+13=0\] send y terms to the other side : \[\large 2x-13= y^2+6y\]

ganeshie8 (ganeshie8):

complete the square for right hand side

ganeshie8 (ganeshie8):

knw how to complete the square ?

OpenStudy (samsan9):

um dont you square root both sides ?

ganeshie8 (ganeshie8):

nope, we're trying to change it into below form : \[\large \color{red}{x-h = a(y-k)^2}\]

OpenStudy (samsan9):

so dont we add the 13 to both sides i am so sorry i honestly had a hard time converting into vertex my last packet :/

ganeshie8 (ganeshie8):

converting to vertex form is easy, all we need to know is the identity : \(\large \color{red}{a^2+2ab+b^2 = (a+b)^2}\)

ganeshie8 (ganeshie8):

\[\large 2x-13= y^2+6y \] \[\large 2x-13= y^2+2(3)y \] add 3^2 both sides : \[\large 2x-13 + \color{green}{3^2}= y^2+2(3)y + \color{green}{3^2}\] \[\large 2x-13 + \color{green}{3^2}= (y+3)^2\]

ganeshie8 (ganeshie8):

simplifying the left hand side gives you : \[\large 2x-4= (y+3)^2\] \[\large 2(x-2)= (y+3)^2\] \[\large x-2= \dfrac{1}{2}(y+3)^2\]

ganeshie8 (ganeshie8):

Now we're ready to compute vertex/focus etc... using this form

ganeshie8 (ganeshie8):

compare above equation with the equation in second column :

ganeshie8 (ganeshie8):

before comparing, lets send the 2 to right side : \[\large x = \dfrac{1}{2}(y+3)^2 + 2\]

ganeshie8 (ganeshie8):

now compare :) vertex = \((h, k )\) = ?

OpenStudy (samsan9):

(2,3)?

OpenStudy (samsan9):

oh yes my bad

ganeshie8 (ganeshie8):

no its my bad ! you're right, I have misinterpreted the equation :o let me fix it quick

ganeshie8 (ganeshie8):

compare below two : \[\large \large x = \dfrac{1}{2}(y+3)^2 + 2 \] \[\large \color{Red}{ x = a(y-k)^2 + h}\]

ganeshie8 (ganeshie8):

Vertex = \(\large (h, k) = (2, -3)\)

ganeshie8 (ganeshie8):

fine ? what else we need to find ?

OpenStudy (samsan9):

the directirx and the length of the focal chord

ganeshie8 (ganeshie8):

lets find the directrix, look at the table and can you tell me the equation for Directrix ? :)

ganeshie8 (ganeshie8):

you need to look at second column in the table..

OpenStudy (samsan9):

y=k-c or x=h-c

ganeshie8 (ganeshie8):

our equation is in the form shown in the last column, so the equation of directrix is \(\large x = h-c\)

ganeshie8 (ganeshie8):

we know \(\large h = 2\) how to find \(\large c\) ?

ganeshie8 (ganeshie8):

use this from the table : \[\large a = \dfrac{1}{4c}\]

ganeshie8 (ganeshie8):

compare the equations again, whats the value of \(a\) ?

OpenStudy (samsan9):

1/2

ganeshie8 (ganeshie8):

yes ! \[\large \dfrac{1}{2} = \dfrac{1}{4c}\] \[\large c = \dfrac{1}{2}\]

ganeshie8 (ganeshie8):

whats the equation of directrix ?

OpenStudy (samsan9):

y=2-1/2

ganeshie8 (ganeshie8):

Yep! but nobody leaves it like that... simplify :)

ganeshie8 (ganeshie8):

hey wait, its `x = 2-1/2` right ?

OpenStudy (samsan9):

oh yeah i got them mixed up

OpenStudy (samsan9):

?

ganeshie8 (ganeshie8):

so what does that simplify to

ganeshie8 (ganeshie8):

directrix : \(\large x = \dfrac{3}{2}\)

OpenStudy (samsan9):

what is the length of the focal chord?

ganeshie8 (ganeshie8):

memorize this formula : length of focal chord = \(\large 4c\)

ganeshie8 (ganeshie8):

in our case \(c = \frac{1}{2}\), so focal chord = ?

OpenStudy (samsan9):

2

ganeshie8 (ganeshie8):

yes ! and u may add above formula to the previous table...

OpenStudy (samsan9):

oh and my bad i just need the focus for it and its done

ganeshie8 (ganeshie8):

oh we still need to find focus is it, its easy, we just look at the last column in table :) whats the focus formula ?

OpenStudy (samsan9):

(h,k+c) or (h+c,k)

ganeshie8 (ganeshie8):

whats there in the right most last column ?

OpenStudy (samsan9):

(h+c,k)?

ganeshie8 (ganeshie8):

Yep !

ganeshie8 (ganeshie8):

do u have all the values ? just plug them

OpenStudy (samsan9):

(5/2,-3)

ganeshie8 (ganeshie8):

Perfect !

ganeshie8 (ganeshie8):

you will need to memorize the table somehow... atleast for the duration you're in conics module

OpenStudy (samsan9):

do you know any good sites to be able to draw a graph for this?

ganeshie8 (ganeshie8):

yes below is best for graphing online https://www.desmos.com/calculator

ganeshie8 (ganeshie8):

enter the equation `y^2+6y-2x+13=0` in left top textbox and press enter

OpenStudy (samsan9):

cool thank you for taking this long with me i really appreciate it and you really helped me alot this summer :)

ganeshie8 (ganeshie8):

good to hear that :) yw :)

OpenStudy (samsan9):

does the site work for any conic sections btw :o

ganeshie8 (ganeshie8):

yes it works for graphing all kinds of equations !

ganeshie8 (ganeshie8):

a bit more sophisticated online graphing tool : http://www.geogebra.org/webstart/geogebra.html

OpenStudy (samsan9):

do you know another website :P?